OpenStudy (anonymous):

how to solve -16t^2+96t+180

OpenStudy (anonymous):

You could use the quadratic equation, to solve the solutions of t (which is what I assume you want) \[(-b \pm \sqrt{b^2-4*a*c})/2*a\]

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

actually what you wrote cannot be solved because it has no equals sign did you leave out the fact that all this equals zero or is your goal to factor it?

OpenStudy (anonymous):

need to determine the maximum height a rock reaches

OpenStudy (lgbasallote):

do you know derivatives?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

aww bummer..it wouldve made this a lot easier

OpenStudy (anonymous):

how owuld you do it with the derrivitives?

OpenStudy (asnaseer):

do you know how to write this in the vertex form? standard form is: \(y=at^2+bt+c\) vertex form is: \(y=a(t-h)^2+k\) where 'h' is the value of 't' at which the vertex occurs

OpenStudy (asnaseer):

if you convert it to vertex form then the max height will occur at t=h and will equal the value of k.

OpenStudy (anonymous):

a rock is propelled up ward from the top of a building 180 ft tall and an inital velocity of 96feet per second. The function that decribes the height of the rocket in the terms of time T is f(t)=-16t^2+96t+180 determine the maximum height the rock reaches.

OpenStudy (anonymous):

maybe this will make more sense with the whole question